# Modeling of AR 12760 with GX Simulator and Evidence for the Extended Transition Region in Peripheral Active Region Loops

### v1

@@ 2.5 Fitting of the Model
The fit is done by comparing on a pixel by pixel basis the SDO/AIA image with the intensity image calculated by GX Simulator. We fit the $240\times 240\arcsec$ region show in Figure 2. Most of the calculated model images had a resolution of 2″ (resulting in a $120\times 120$ pixel field of view), and were convolved with a 1.2″ point spread function, but, as will be discussed in Section 3.1, we also did fits at 8″ resolution. The AIA data were rescaled to match.
@@
GX Simulator does any needed alignment of the AIA maps with the magnetic field model (see Section 2.3). Unlike the radio modeling done with GX Simulator (e.g., Fleishman et al., 2021b), the relative pointings of the data from the two SDO instruments is well understood, so no cross-correlation based alignment was done.
@@
When calculating the model image we constrain the transition region emission to areas where $B_{z}\geq 200$ G in the magnetogram, as in Nita et al. (2018). The resulting transition region mask is shown in Figure 4. This confines the transition region emission to those areas where the photospheric field is strong and mostly vertical (Klimchuk, 1987; Nita et al., 2018). The mask accounts approximately for the rapid expansion of the field in the upper photosphere and chromosphere, which occurs because of the transition from very large to very small plasma beta. This transition occurs over a thin but finite layer. In contrast, the magnetic field extrapolation assumes an infinitely sharp boundary, which leads to strong transition region emission everywhere, contrary to observations.
@@
To emphasize the bright regions of the map, while still incorporating constrains from the the dark areas surrounding the active region, we assigned a constant uncertainty to all pixels in a given waveband, so that our model-to-data comparison metric is the sum of the squared residuals between the model and data, $\sum\delta^{2}=\sum(I_{\mbox{\scriptsize obs}}-I_{\mbox{\scriptsize mod}})^{2}$, where $I_{\mbox{\scriptsize obs}}$ and $I_{\mbox{\scriptsize mod}}$ are the observed and modeled intensities respectively. This means that the brighter areas of the loop are weighted more heavily, but the emission is still constrained to low values outside the active region. This metric of success allows us to compare the quality of the fit within a given waveband, but it does not allow us to provide quantitative goodness of fit contours in our parameter space, or compare the goodness of fits between wavebands.
@@
For each ($a,b$) pair, we employed the Coronal Heating Modeling Pipeline utility (CHMP, Nita et al., 2023), which is part of the GX Simulator package, to compute maps for a range of values of $q_{0}$ converging to a minimum $\sum\delta^{2}$ value, and thus determining the best fit for that set of ($a,b$) values. We considered a range of values for $a$ from 1-1 to 3 and $b$ from 1-1 to 2.5, in increments of 0.5.
@@
We did two general types of data-model comparisons. Most included both the modeled corona and transition region emission. As mentioned above, the modeled transition region emission is limited to the regions with photospheric $B_{z}\geq 200$ G as shown in Figure 4. All points in the array were used in the data-model comparison except for a one pixel wide band around the edge of the array to allow for shifts of the mask when interpolating to align with the AIA data. We also did fits modeling coronal emission only. In these fits the model transition region emission was set to zero and the $\sum\delta^{2}$ calculation excluded the points designated as transition region as well as the edge pixels.
@@
Below we focus on two AIA bands, 211 Å which shows chiefly coronal emission, and 171 Å which shows a combination of the coronal and transition region emission.
3 Fit Results
3.1 Fit to the 211 Å band
@@
Figure 5: Thumbnails of model 211 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for Magnetic Model 1, a model/data comparison resolution of 2″, and including the model transition region. The actual 211 Å image at the same resolution is at the bottom left. The $a=1$, $b=1$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. Maps with $\sum\delta^{2}$  within 10% and 20% of the minimum have titles and borders in orange and red respectively.
@@
We first present the fits to the 211 Å band, which is dominated by Fe XIV with a peak emission temperature $\log T\mbox{(K)}\approx 6.3$ (Boerner et al., 2012). In Figure 5 we show a grid of thumbnails showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for Model 1 including emission contributed by both the model corona and transition region. The best fit corresponds to $a=1$, $b=1$, $q_{0}=4.4\times 10^{-3}$ erg cm-3 s-1. This model image has a title and border in green. Maps with $\sum\delta^{2}$  within 10% and 20% of the minimum have titles and borders in orange and red respectively. The result shows a linear area with similar fit quality. We discuss this in more detail in Section 4.1, which includes a plot of $\sum\delta^{2}$.
@@
The best fits for the 2″ and 8″ comparisons are shown in Figure 6 and Figure 7 respectively, along with the AIA data at the same resolution, and the residual, $\delta=I_{obs}-I_{mod}$. The plots also show the model coronal and transition region contributions to the emission as separate maps. The model maps look reasonable on a general level, with most of the emission coming from the active region core. However, the detailed correspondence of the bright features is not as good as we would like. A significant part of this can be attributed to inaccuracies in the magnetic field extrapolation (Fleishman et al., 2019). Even if the hydrodynamics along field lines is modeled correctly, the spatial location of the emission will be off. We mitigate this in part by reducing the resolution to 8″, although differences are still apparent. The transition region emission is a significant contributor to the emission near loop footpoints, but is not the main source of the 211 Å emission.
@@
Figure 6: Comparison of the best fit to 211 Å data binned to 2″ resolution showing (a) the AIA data, (b) the best fit model map, (c) The residual, $\delta=I_{obs}-I_{mod}$, (d) the coronal component of the model, and (e) the transition region component of the model. Panels (a),(b),(d), and (e) are scaled linearly from 0 to 95% of the maximum data intensity. Panel (c) is scaled to $\pm 95\%$ of the maximum data intensity.
@@
Figure 7: Comparison of the best fit to 211 Å data binned to 8″ resolution showing (a) the AIA data, (b) the best fit model map, (c) The residual, $\delta=I_{obs}-I_{mod}$, (d) the coronal component of the model, and (e) the transition region component of the model. Panels (a),(b),(d), and (e) are scaled linearly from 0 to 95% of the maximum data intensity. Panel (c) is scaled to $\pm 95\%$ of the maximum data intensity. The regions in green in the residual map were not used to calculate $\sum\delta^{2}$.
@@
The transition region emission in the model seems especially bright and localized compared to the observations. We address the localization later. The brightness is a problem that has historically afflicted many models, including detailed 1D hydrodynamics models, i.e., “loop” models. As mentioned in Section 2.4, the discrepancy is reduced in expanding loops, but the excess brightness remains an unsolved problem. For this reason, we also considered a fit to the corona only. The resulting best fit model is shown in Figure 8. For this fit we turned off the transition region emission and used only pixels not designated as including transition region (Figure 4) to calculate $\sum\delta^{2}$. In this case we found the best fit to be $a=2$, $b=0.5$, $q_{0}=5.8\times 10^{-3}$ erg cm-3 s-1. The intensity scale is the same as in Figure 6. The model looks similar to the best fit model for the combined corona and transition region fit, although is a bit brighter than the corona alone in that fit. This ($a,b$) pair also corresponds to a pair that is within 10% of the minimum $\sum\delta^{2}$ for the corona and transition region fit shown in Figure 5.
@@
Figure 8: AIA data (a) and best fit (b) to the corona only (leaving out transition region emission) and the residual values ($\delta=I_{\mbox{\scriptsize obs}}-I_{\mbox{\scriptsize mod}}$) for 211 Å using Magnetic Model 1 and a fit comparison resolution of 2″. The regions in green in the residual map were not used to calculate $\sum\delta^{2}$. These include the areas we expected to contain transition region emission (See Figure 4).
3.2 Fit to the 171 Å band
@@
Figure 9: Thumbnails of model 171 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for Model 1, a model/data comparison resolution of 2″, and including both the corona and transition region. The actual 171 Å image at the same resolution is at the bottom left. The $a=0.5$, $b=0.0$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. Maps with $\sum\delta^{2}$  within 10% and 20% of the minimum have titles and borders in orange and red respectively.
The difference appears to be connected to longer loops on the periphery of the active region. Figure 10 shows the details of the best fit model, again with separate images showing the corona and transition region components. As would be expected, the localized model transition region emission is more evident in the 171 Å band than in 211 Å. However, in the “best fit” this seems to be exaggerated; almost all the emission is coming from the designated transition region foot points with very little emission in the active region core. Also, although the 171 Å image is dominated by the bright areas in the long, peripheral loops (circled in red in Figure 10a), these features are entirely missing in the model images. This is not just true of the best fit, but of all the fits shown in Figure 9. Feature A appears to be the base of a fan-loop, a large scale structure that we think closes in the corona. Feature B seems to form a more (although not entirely) complete arc, although its exact size and altitude are not clear.
Figure 10: Comparison of the best fit to 171 Å data binned to 2″ resolution and using Magnetic Model 1 showing (a) the AIA data, (b) the best fit model map, (c) The residual, $\delta=I_{\mbox{\scriptsize obs}}-I_{\mbox{\scriptsize mod}}$, (d) the coronal component of the model, and (e) the transition region component of the model. Panels (a),(b),(d), and (e) are scaled linearly from 0 to 95% of the maximum data intensity. Panel (c) is scaled from $\pm 95\%$ of the maximum data intensity.
One possible explanation for the total lack of emission in the peripheral loops is that the relevant loops did not close within the modeled cube, and thus were not included in the emission (see Section 2.3). This was the reason we developed Magnetic Model 2. Figure 11 shows magnetic field lines from both Magnetic Model 1 and Model 2 superimposed over the 171 Å reference image. It can be seen that in Magnetic Model 1 some of the longer field lines extending from the periphery of the active region do not close down again inside of the box (these are shown in yellow), and so are being excluded from our model emission. In contrast, in Magnetic Model 2, the lines from these areas do close down inside the NLFFF model cube, so we would expect them to be included in the emission.
@@
However, this proposed explanation does not seem to be correct. In Figure 12 we show the best fit result for model maps calculated with Magnetic Model 2. The best fit is $a=1.0$, $b=0.0$, $q_{0}=4.7\times 10^{-4}$ erg cm-3 s-1. The full ($a,b$) grid is shown in the Appendix, in Figure 19. The longer field lines in the periphery of the active region are still not present in the emission, although the footpoints of the loops in the transition region are very bright.
Thus, the use of the lower-resolution but larger-volume magnetic field model results in a less accurate distribution of emission in the core of the active region, but it does not result in the appearance of the longer loops, so the problem does not seem to be that the loops are being left out of the magnetic field model. As we will discuss in Section 4.2, we think the problem is in how the model transition region emission is distributed along the loop.
Figure 12: Comparison of the best fit to 171 Å data using Magnetic Model 2 and binned to 2″ resolution showing (a) the AIA data (b) the best fit model map, (c) The residual, $\delta=I_{\mbox{\scriptsize obs}}-I_{\mbox{\scriptsize mod}}$, (d) the coronal component of the model, and (e) the transition region component of the model. Panels (a),(b),(d), and (e) are scaled linearly from 0 to 95% of the maximum data intensity. Panel (c) is scaled to $\pm 95\%$ of the maximum data intensity.
3.3 Fits to other AIA bands
The other AIA bands we considered, 94, 131, 193, and 335 Å, show a temperature dependent pattern similar to that seen in 171 and 211 Å. For the 211 and 335 Å  bands, which are dominated by emission at $\log T\mbox{(K)}\gtrsim 6.3$ (Boerner et al., 2012), the fits look similar. Although they differ from the original AIA image in detail, they show similar amount of emission in the active region core, which dominates the emission in both data and model. The thumbnail plots for the comparisons of the model to the 335 Å data binned to 8″ are shown in Figure 20, and more discussion of the 211 Å and 335 Å results is in Section 4.1.

### v2

@@ II.5 Fitting of the Model
The fit is done by comparing on a pixel by pixel basis the SDO/AIA image with the intensity image calculated by GX Simulator. We fit the $240\times 240\arcsec$ region show in Figure 2. Most of the calculated model images had a resolution of 2″ (resulting in a $120\times 120$ pixel field of view), and were convolved with a 1.2″ point spread function, but, as will be discussed in Section III.1, we also did fits at 8″ resolution. The AIA data were rescaled to match, with the intensities rescaled to account for the change in the pixel area.
@@
GX Simulator does any needed alignment of the AIA maps with the magnetic field model (see Section II.3). Unlike the radio modeling done with GX Simulator (e.g., Fleishman et al., 2021b), the relative pointings of the data from the two SDO instruments is well understood, so no cross-correlation based alignment was done.
@@
When calculating the model image we constrain the transition-region emission to areas where $|B_{z}|\geq 200$ G in the magnetogram, as in Nita et al. (2018). The resulting transition-region mask is shown in Figure 5. As stated in the pervious section, all transition-region emission is restricted to the lowest layer of the model data cube. With this additional mask, we also confine the transition-region emission to only those areas where the photospheric field is strong and mostly vertical (Klimchuk, 1987; Nita et al., 2018). The mask accounts approximately for the rapid expansion of the field in the upper photosphere and chromosphere, which occurs because of the transition from very large to very small plasma beta. This transition occurs over a thin but finite layer. In contrast, the magnetic field extrapolation assumes an infinitely sharp boundary, which leads to unrealistically high magnetic flux in areas with lower magnetic field, and thus strong transition-region emission everywhere, contrary to observations. Thus we suppress the model transition-region emission in these low $B_{z}$ regions. The physical motivation for the mask is discussed further by Nita et al. (2018) and in Appendix A.
@@
To emphasize the bright regions of the map, while still incorporating constrains from the the dark areas surrounding the active region, we assigned a constant uncertainty to all pixels in a given wave band, so that our model-to-data comparison metric is the sum of the squared residuals between the model and data, $\sum\delta^{2}=\sum(I_{\mbox{\scriptsize obs}}-I_{\mbox{\scriptsize mod}})^{2}$, where $I_{\mbox{\scriptsize obs}}$ and $I_{\mbox{\scriptsize mod}}$ are the observed and modeled intensities respectively. This means that the brighter areas of the image are weighted more heavily, but the emission is still constrained to low values outside the active region. This metric of success allows us to compare the quality of the fit within a given wave band, but it does not allow us to provide quantitative goodness-of-fit contours in our parameter space, or compare the goodness of fits between wave bands.
@@
For each ($a,b$) pair, we employed the Coronal Heating Modeling Pipeline utility (CHMP, Nita et al., 2023), which is part of the GX Simulator package, to compute maps for a range of values of $q_{0}$ converging to a minimum $\sum\delta^{2}$ value, thus determining the best fit for that set of ($a,b$) values. For areas of the ($a,b$) grid with the lower $\sum\delta^{2}$ values the $q_{0}$ vs. $\sum\delta^{2}$ curves were smooth with clear minima. We considered a range of values for $a$ from 1-1 to 3 and $b$ from 1-1 to 2.5, in increments of 0.5.
@@
We did two general types of data-model comparisons. Most included both the modeled corona and transition-region emission. As mentioned above, the modeled transition-region emission is limited to the bottom layer of the model cube in regions with photospheric $|B_{z}|\geq 200$ G as shown in Figure 5. All points in the array were used in the data-model comparison except for a one pixel wide band around the edge of the array to allow for shifts of the mask when interpolating to align with the AIA data. We also did fits modeling coronal emission only. GX Simulator calculates the corona and transition region separately (as discussed in Section II.4) and the transition-region emission can be turned off entirely. Thus in the corona-only fit the model transition-region emission was set to zero and the $\sum\delta^{2}$ calculation excluded the points designated as transition region (Figure 5) as well as the edge pixels.
@@
Below we focus on two AIA bands, 211 Å, which shows chiefly coronal emission, and 171 Å, which shows a combination of coronal and transition-region emission.
III Fit Results
III.1 Fit to the 211 Å band
@@
Figure 6: Thumbnails of model 211 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for Magnetic Model 1, with a model/data comparison resolution of 2″, and including the model transition region. The actual 211 Å image at the same resolution is at the bottom left. The $a=1$, $b=1$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. See Fig. 7b for a larger version of the best model. Maps with $\sum\delta^{2}$ within 10% and 20% of the minimum have titles and borders in orange and red, respectively. The images are all scaled linearly with the same intensity range.
@@
We first present the fits to the 211 Å band, which is dominated by Fe XIV with a peak emission temperature $\log T\mbox{(K)}\approx 6.3$ (Boerner et al., 2012). In Figure 6 we show a grid of thumbnails showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for Model 1, including emission contributed by both the model corona and transition region. The best fit corresponds to $a=1$, $b=1$, $q_{0}=4.4\times 10^{-3}$ erg cm-3 s-1. This model image has a title and border in green. Maps with $\sum\delta^{2}$ within 10% and 20% of the minimum have titles and borders in orange and red, respectively. The result shows a linear area with similar fit quality. We discuss this in more detail in Section IV.1, which includes a plot of $\sum\delta^{2}$.
@@
The best fits for the 2″ and 8″ comparisons are shown in Figure 7 and Figure 8 respectively, along with the AIA data at the same resolution, and the residual, $\delta=I_{obs}-I_{mod}$. The plots also show the model coronal and transition-region contributions to the emission as separate maps. The coronal and transition-region emissions are calculated separately and summed by GX Simulator, and seperate transition-region and coronal maps can be synthesized in addition to the total emission. The model maps look reasonable on a general level, with most of the emission coming from the active region core. However, the detailed correspondence of the bright features is not as good as we would like. A significant part of this can be attributed to inaccuracies in the magnetic field extrapolation (Fleishman et al., 2019). Even if the hydrodynamics along field lines is modeled correctly, the spatial location of the emission will be off. We mitigate this in part by reducing the resolution to 8″, although differences are still apparent. The transition-region emission is a significant contributor to the emission near loop footpoints, but it is not the main source of the 211 Å emission.
@@
Figure 7: Comparison of the best fit to 211 Å data binned to 2″ resolution using Magnetic Model 1 showing (a) the AIA data, (b) the best-fit model map, (c) The residual, $\delta=I_{obs}-I_{mod}$, (d) the coronal component of the model, and (e) the transition-region component of the model. Panels (a),(b),(d), and (e) are scaled linearly from 0 to 95% of the maximum data intensity. Panel (c) is scaled to $\pm 95\%$ of the maximum data intensity. See Fig. 6 for grid of all ($a,b$) best matches.
@@
Figure 8: Comparison of the best fit to 211 Å data binned to 8″ resolution using Magnetic Model 1 showing (a) the AIA data, (b) the best-fit model map, (c) The residual, $\delta=I_{obs}-I_{mod}$, (d) the coronal component of the model, and (e) the transition-region component of the model. Panels (a),(b),(d), and (e) are scaled linearly from 0 to 95% of the maximum data intensity. Panel (c) is scaled to $\pm 95\%$ of the maximum data intensity. The regions in green in the residual map were not used to calculate $\sum\delta^{2}$. See Fig. 18 for grid of all ($a,b$) best matches.
@@
The transition-region emission in the model seems especially bright and localized compared to the observations. We address the localization later. The brightness is a problem that has historically afflicted many models, including detailed 1D hydrodynamics models, i.e., “loop” models. As mentioned in Section II.4, the discrepancy is reduced in expanding loops, but the excess brightness remains an unsolved problem. For this reason, we also considered a fit to the corona only. The resulting best-fit model is shown in Figure 9. For this fit we turned off the transition-region emission and used only pixels not designated as including transition region (Figure 5) to calculate $\sum\delta^{2}$. In this case we found the best fit to be $a=2$, $b=0.5$, $q_{0}=5.8\times 10^{-3}$ erg cm-3 s-1. The intensity scale is the same as in Figure 7. The model looks similar to the best-fit model for the combined corona and transition-region fit, although is a bit brighter than the corona alone in that fit. This ($a,b$) pair also corresponds to a pair that is within 10% of the minimum $\sum\delta^{2}$ for the corona and transition-region fit shown in Figure 6.
@@
Figure 9: AIA data (a) and best fit (b) to the corona only (leaving out transition-region emission) and the residual values ($\delta=I_{\mbox{\scriptsize obs}}-I_{\mbox{\scriptsize mod}}$) for 211 Å using Magnetic Model 1 and a fit comparison resolution of 2″. The regions in green in the residual map were not used to calculate $\sum\delta^{2}$. These include the areas we expected to contain transition-region emission (See Figure 5). See Fig. 19 for grid of all ($a,b$) best matches.
III.2 Fit to the 171 Å band
@@
Figure 10: Thumbnails of model 171 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for Model 1, a model/data comparison resolution of 2″, and including both the corona and transition region. The actual 171 Å image at the same resolution is at the bottom left. The $a=0.5$, $b=0.0$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. See Fig. 11b for larger version of the best model. Maps with $\sum\delta^{2}$ within 10% and 20% of the minimum have titles and borders in orange and red, respectively. The images are all scaled linearly with the same intensity range.
The difference appears to be connected to longer loops on the periphery of the active region. Figure 11 shows the details of the best-fit model, again with separate images showing the corona and transition-region components. As would be expected, the localized model transition-region emission is more evident in the 171 Å band than in 211 Å. However, in the “best fit” this seems to be exaggerated; almost all the emission is coming from the designated transition-region foot points with very little emission in the active region core. Also, although the 171 Å image is dominated by the bright areas in the long, peripheral loops (circled in red in Figure 11a), these features are entirely missing from the model images. This is not just true of the best fit, but of all the fits shown in Figure 10. Feature A appears to be the base of a fan-loop, a large scale structure that we think closes in the corona. Feature B seems to form a more (although not entirely) complete arc, although its exact size and altitude are not clear.
Figure 11: Comparison of the best fit to 171 Å data binned to 2″ resolution and using Magnetic Model 1 showing (a) the AIA data, (b) the best-fit model map, (c) The residual, $\delta=I_{\mbox{\scriptsize obs}}-I_{\mbox{\scriptsize mod}}$, (d) the coronal component of the model, and (e) the transition-region component of the model. Panels (a),(b),(d), and (e) are scaled linearly from 0 to 95% of the maximum data intensity. Panel (c) is scaled from $\pm 95\%$ of the maximum data intensity. See Fig. 10 for grid of all ($a,b$) best matches.
One possible explanation for the total lack of emission in the peripheral loops is that the relevant loops did not close within the modeled cube, and thus were not included in the emission (see Section II.3). This was the reason we developed Magnetic Model 2. Figure 12 shows magnetic field lines from both Magnetic Model 1 and Model 2 superimposed over the 171 Å reference image. It can be seen that in Magnetic Model 1 some of the longer field lines extending from the periphery of the active region do not close down again inside of the box (these are shown in yellow), and so are being excluded from our model emission. In contrast, in Magnetic Model 2, the lines from these areas do close down inside the NLFFF model cube, so we would expect them to be included in the emission.
@@
However, this proposed explanation does not seem to be correct. In Figure 13 we show the best fit result for model maps calculated with Magnetic Model 2. The best fit is $a=1.0$, $b=0.0$, $q_{0}=4.7\times 10^{-4}$ rg cm-3 s-1. The full ($a,b$) grid is shown in Appendix B, in Figure 20. The longer field lines in the periphery of the active region are still not present in the emission, although the footpoints of the loops in the transition region are very bright.
Thus, the use of the lower-resolution but larger-volume magnetic field model results in a less accurate distribution of emission in the core of the active region, but it does not result in the appearance of the longer loops, so the problem does not seem to be that the loops are being left out of the magnetic field model. As we will discuss in Section IV.2, we think the problem is in how the model transition-region emission is distributed along the loop.
Figure 13: Comparison of the best fit to 171 Å data using Magnetic Model 2 and binned to 2″ resolution showing (a) the AIA data (b) the best-fit model map, (c) The residual, $\delta=I_{\mbox{\scriptsize obs}}-I_{\mbox{\scriptsize mod}}$, (d) the coronal component of the model, and (e) the transition-region component of the model. Panels (a),(b),(d), and (e) are scaled linearly from 0 to 95% of the maximum data intensity. Panel (c) is scaled to $\pm 95\%$ of the maximum data intensity. See Fig. 20 for grid of all ($a,b$) best matches.
III.3 Fits to other AIA bands
The other AIA bands we considered, 94, 131, 193, and 335 Å, show a temperature dependent pattern similar to that seen in 171 and 211 Å(see Figure 14). For the 211 and 335 Å bands, which are dominated by emission at $\log T\mbox{(K)}\gtrsim 6.3$ (Boerner et al., 2012), the fits look similar. Although they differ from the original AIA image in detail, they show similar amount of emission in the active region core, which dominates the emission in both data and model. The thumbnail plots for the comparisons of the model to the 335 Å data binned to 8″ are shown in Figure 21, and more discussion of the 211 Å and 335 Å results is in Section IV.1.